OpenStudy (anonymous):

divide 3x^3+10x^2-9x-4 by x+4

OpenStudy (anonymous):

\[\left(\begin{matrix}x+4 \\3x ^{3}+10x ^{2}-9x -4\end{matrix}\right)=\left(\begin{matrix}x +4 \\ \left( x -1\right)\ \left(3x +1 \right)\left( x +4 \right) \end{matrix}\right)=\left(\begin{matrix}1 \\ \left(x -1 \right)\left( 3x +1 \right)\end{matrix}\right)\]

OpenStudy (anonymous):

\[3 x^2-2 x-1 \] with no remainder.